OpenStudy (anonymous):

If the equations x^2+px+q=0 and x^2+bx+c=0 have 1 common root, then it is equal to?? Ans.-[(q-c)/(p-b)]

OpenStudy (anonymous):

x^2+px+q-(x^2+bx+c)=px+q-bx-c=0 x=-[(q-c)/(p-b)]

OpenStudy (shubhamsrg):

let a be common root.. then: a^2 +pa + q =0 and a^2 +ab +c=0 subtracting both : you get a = your ans